FOURTH HUNGARIAN COLLOQUIUM ON LIMIT THEOREMS

OF PROBABILITY AND STATISTICS Balatonlelle, June 28– July 2, 1999

Scientific Program

Monday, June 28 Morning

10.00–10.30  Opening

10.30–11.00 Coffee break

11.00–11.45  M. Csörgő
11.45–12.30  P. Deheuvels

Afternoon

Section A  Section B
2.00–2.25  L. Rejtő  D. de Waal
2.25–2.50  A. Nádas  T. Lin
2.50–3.15  A. Járai  J. Beirlant
3.15–3.40  A. Jurlewicz  G. Draisma

3.40–4.00 Coffee break

4.00–4.25  A. Zaihraiev  H. Liero
4.25–4.50  G. L. O’Brien  D. M. Dabrowska
4.50–5.15  D. Petz  A. Antos
5.15–5.40  E. Ursianu  I. Van Keilegom
5.40–6.05  S. Baran

19.30 Conference reception Tuesday, June 29 Morning

8.30– 9.00  R. Bhattacharya
9.00– 9.30  S. J. Taylor
9.30–10.00  M. Yor
10.00–10.30  Yu. Davydov

10.30–11.00 Coffee break

11.00–11.30  W. Philipp
11.30–12.00  R. Bradley
12.00–12.30  M. R. Leadbetter

Afternoon

Section A  Section B
2.00–2.25  M. Hušková  A. Zaitsev
2.25–2.50  J. Steinebach  K. Fukuyama
2.50–3.15  E. Gombay  B. Courbot
3.15–3.40  M. Orasch  Ch. Suquet

3.40–4.00 Coffee break

4.00–4.25  T. Inglot  A. Spătaru
4.25–4.50  M. D. Burke  S. Borovkova
4.50–5.15  M. Bloznelis  M. Janžura
5.15–5.40  A. Gnedin  I. Fazekas
5.40–6.05  G. Pritchard

Wednesday, June 30 Morning

Section A  Section B
9.00– 9.25  M. Lifshits  P. Lachout
9.25– 9.50  P. Major  Z. S. Szewczak
9.50–10.15  I. Fahrner  E. Liebscher

10.30–11.00 Coffee break

10.35–11.00  G. Simons  P. Szabłowski
11.00–11.25  Z. Megyesi  S. Ostrovska
11.25–11.50  G. Christoph  L. Ericksen
11.50–12.15  A. V. Nagaev  A. Plucinska

Afternoon Excursion Thursday, July 1 Morning

Section A  Section B
9.00– 9.25  I. Berkes  D. Hlubinka
9.25– 9.50  J. Mielniczuk  L. Gerencsér
9.50–10.15  P. Mörters  V. A. Dikarev

10.30–11.00 Coffee break

10.35–11.00  I. Csiszár  L. Rüschendorf
11.00–11.25  D. M. Chibisov  G. Morvai
11.25–11.50  L. Horváth  E. Dyakonova
11.50–12.15  F. Lombard  V. Kolchin

Afternoon

Section A  Section B
2.00–2.25  L. Devroye
2.25–2.50  S. Csörgő  A. Krzyżak
2.50–3.15  T. Móri  V. Zaiats
3.15–3.40  G. Haiman  S. Döhler

3.40–4.00 Coffee break

4.00–4.25  K. Yamada  J. Lember
4.25–4.50  V. de la Peña  L. Viharos
4.50–5.15  R. Zitikis  N. Mushkudiani
5.15–5.40  Y. Hu  M. Horváth
5.40–6.05  Ch. El-Nouty  Gy. Terdik
6.05–6.30  T. Szabados

Friday, July 2 Morning

8.35– 9.00  H. Kaspi
9.00– 9.25  Z. Shi
9.25– 9.50  A. Lachal
9.50–10.15  Ph. Berthet

10.30–11.00 Coffee break

10.35–11.00  B. Gittenberger
11.00–11.25  N. Eisenbaum
11.25–11.50  A. Földes
11.50–12.15  E. Csáki

Poster session: Tuesday, June 29, 6.05 pm.